OpenStudy (anonymous):

Linear Algebra:

OpenStudy (anonymous):

For all integers n, \[n^3-n\] is divisible by 3. How to show this?

OpenStudy (xapproachesinfinity):

what did you try so far?

OpenStudy (anonymous):

I know that I can use induction method but not sure how to apply it yet so i've been trying to manipulate it to prove it by that method

OpenStudy (loser66):

Hint: factor n out, simplify to get the logic

OpenStudy (anonymous):

\[n(n^2-1)\]

OpenStudy (anonymous):

This is what i tried also but not sure how that helps me?

OpenStudy (anonymous):

to show it is divisible by 3

OpenStudy (loser66):

more

OpenStudy (xapproachesinfinity):

n(n-1)(n+1)

OpenStudy (loser66):

Yup

OpenStudy (anonymous):

Would I then try to get a 1/3 out?

OpenStudy (xapproachesinfinity):

Actually this not true for any integer?

OpenStudy (xapproachesinfinity):

you can try out some integers

OpenStudy (anonymous):

\[(2)(2-1)(2+1) = 2(1)(3)\]

OpenStudy (anonymous):

So i should set the condition: \[n \ge 1\]

OpenStudy (xapproachesinfinity):

you goal is to prove that n(n-1)(n+1)=3k that is what it means to be divisible by 3

OpenStudy (xapproachesinfinity):

what about negative intergers? what did the original problem state

OpenStudy (anonymous):

for every positive integer is only thing that was stated

OpenStudy (xapproachesinfinity):

okay good you stated for any integer n, which made think that there must be a problem

OpenStudy (xapproachesinfinity):

if it is positive integers that's good!

OpenStudy (anonymous):

But when we use 1 there is a zero ? \[(1)(0)(2) = 0\]

OpenStudy (xapproachesinfinity):

well 0 divisible by any number we get back 0

OpenStudy (anonymous):

True so then k is also zero

OpenStudy (loser66):

hey, Positive integer means >0 (strictly )

OpenStudy (loser66):

Non-negative is \(\geq \)0

OpenStudy (xapproachesinfinity):

eh good catch @Loser66 zero is not included

OpenStudy (xapproachesinfinity):

my bad!

OpenStudy (anonymous):

so zero if zero is not included there must still be a problem?

OpenStudy (xapproachesinfinity):

what problem?

OpenStudy (xapproachesinfinity):

no you got me wrong! i meant not included in positive integers but 0 is still divisible by 3

OpenStudy (anonymous):

Ohhhh okay I see now so by expanding my original expression and plugging in positive integers I showed that it is divisible by 3 but how would i know to add the "3k" to the right of my expression for next time?

OpenStudy (xapproachesinfinity):

not sure i got you point here hehe

OpenStudy (xapproachesinfinity):

i was replying to @Johnbc

OpenStudy (xapproachesinfinity):

yes those consecutive

OpenStudy (xapproachesinfinity):

No worries

OpenStudy (xapproachesinfinity):

eh i have to go lol

OpenStudy (anonymous):

Thank you very much I am still unsure about the right side but I understand the left

OpenStudy (xapproachesinfinity):

good luck with this^_^ @Loser66 gave you a hint to go with this

OpenStudy (anonymous):

Plugging in P(n)?

OpenStudy (anonymous):

Thank you

OpenStudy (xapproachesinfinity):

you can easily see for n=3 as he showed but you need to do it for n and n+1

OpenStudy (xapproachesinfinity):

you are welcome